OpenStudy (anonymous):

Solve for the interval 2sin2x-1=0 interval[0,2pi)

OpenStudy (turingtest):

where are you stuck?

OpenStudy (anonymous):

I found out the answer and it is {pi/12, 5pi/12, 13pi/12, and 17pi/12} I do not understand how to get the 13pi/12 and 17pi/12 !

OpenStudy (turingtest):

ok, so you get to sin2x=1/2 2x=????

OpenStudy (anonymous):

that equals 1

OpenStudy (turingtest):

no... 2x=arcsin(1/2) what is the inverse sine of 1/2 ?

OpenStudy (turingtest):

i.e. at what angle is the sine 1/2 ?

OpenStudy (anonymous):

30 degrees, or \[\pi/6, 5\pi/6\]

OpenStudy (turingtest):

great, so we have\[2x=\{\frac\pi 6,\frac{5\pi} 6\}\] but any time we add \(2\pi\) we will also have the same point on the unit circle, so we can write\[2x=\{\frac\pi 6+2\pi n,\frac{5\pi} 6+2\pi n\};n \in\mathbb N\]

OpenStudy (turingtest):

so far so good?

OpenStudy (anonymous):

yep!

OpenStudy (turingtest):

now we solve for x\[x=\{\frac\pi{12}+\pi n,\frac{5\pi}{12}+\pi n\}; n\in\mathbb N\]do you see how the other answers come up now?

OpenStudy (anonymous):

so I just pick any number, n, and plug it into that equation. I just tried to use n=2 and got \[25\pi/12\]

OpenStudy (turingtest):

ah, but you have missed the fact that when we divided by 2, the +2pi*n became +pi*n plug in n=0 and n=1 now and see what you get

OpenStudy (turingtest):

if you use n=2 that takes us out of the range of 0 to 2pi, so that is invalid here

OpenStudy (turingtest):

you can't just use ANY n to get all answers.... I think I misunderstood the problem you were having

OpenStudy (anonymous):

n=0, ='s \[\pi/12\] and n=1, ='s \[13\pi/12\]

OpenStudy (anonymous):

do I not try to make n=3 or n=4 because the interval only includes 2 pi?

OpenStudy (turingtest):

yes, you have to check how many values of n (n=0,1,2,3) can fit until you leave the range 0<x<2pi

OpenStudy (turingtest):

in this case only n=0 and n=1 give angles in the range specified

OpenStudy (anonymous):

and how do I tell is something is more than 2pi?

OpenStudy (turingtest):

well with n=0 your solutions are pi/12 and 5pi/2 with n=1 they are 13pi/12 and 17pi/12 with n=2 they are 25pi/12 and 29pi/12, both of which are greater than 2pi (since 25/12>2)

OpenStudy (anonymous):

great thank you sooooo much this was very helpful :)

OpenStudy (turingtest):

Glad I could help :) See ya around!